OpenStudy (anonymous):

A company offers a starting salary of $33,000 with raises of $3,500 per year. Find the total salary over a 10 year period.

OpenStudy (anonymous):

you up for a word problem ? @e.mccormick

OpenStudy (anonymous):

What you want to find is 33000 + (33000+3500) + (33000+3500*2) + (33000+3500*3) + ... + (33000+3500*9) If you learned Arithmetic Sequence Formula, starting term is 33000 difference between terms is 3500 there are 10 terms So, the answer would be \[\frac{n(2a+(n-1)d)}{2}=\frac{10(66000+9*3500)}{2}=487500\]

OpenStudy (e.mccormick):

@angelina22309 Hey there. I have been going over area formulas with determinants... so Idid not see your message. How are you doing on this one? Did what nightwill put up make sense?

OpenStudy (anonymous):

It did make sense but the answer is actually $48,750

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

wait!

OpenStudy (anonymous):

Nevermind I forgot the 10 so yea hes right

OpenStudy (e.mccormick):

Yah, the total salary would be what you were paid over all 10 years. The salary AFTER 10 years is a different question.